X+3y=10
-3x+y=-10

Answer by Cromlix(4381)

Hi there,
I will do it by elimination method.
x+3y=10........(1)
-3x+y=-10......(2)
Multiply (1) by 3
3x + 9y = 30......(1)
-3x + y = -10.....(2)
Add (1) and (2)
......10y = 20
........y = 2
Substitute y = 2 into (1)
x + 3y = 10
x + 3(2) = 10
x + 6 = 10
x = 10 - 6
x = 4
{4,2}
Hope this helps :-)